OpenStudy (anonymous):

(2, -3) and (k, 7); m = -2 Find the value of K so that the line through the given points has the given slope.

OpenStudy (zzr0ck3r):

\(\frac{7-(-3)}{k-2}=-2\) can you solve for \(k\)?

OpenStudy (triciaal):

Point slope formula (y2 -y1) = m(x2 - x1) for points (x1 , y1) and (x2, y2) and slope m